OpenStudy (benlindquist):

f(x) = 1/2 x+4 then find the value of f -1^(4)

OpenStudy (shadowlegendx):

Inverse the function, then apply 4 for x

OpenStudy (solomonzelman):

You can just plug in f(x)=4.

OpenStudy (triciaal):

if f(x) =y then switch x and y and solve for y set = 4

OpenStudy (solomonzelman):

it is shorter to just plug in f(x)=4 (then your x is f^(-1)(4).)

OpenStudy (benlindquist):

so fx = 4 then find the value f -1^(4)

OpenStudy (benlindquist):

f -1^(x) blank x blank it has that in the first row

OpenStudy (benlindquist):

f -1^(4)

OpenStudy (solomonzelman):

Example: \(\color{blue}{\displaystyle f(x)=4x+8}\), find \(\color{blue}{\displaystyle f^{-1}(8)}\). -------------------------------------- Official approach. \(\color{blue}{\displaystyle y=4x+8}\) \(\color{blue}{\displaystyle x=4y+8}\) \(\color{blue}{\displaystyle x-8=4y}\) \(\\[0.5em]\) \(\color{blue}{\displaystyle \frac{x-8}{4}=y}\) \(\\[1.1em]\) \(\color{blue}{\displaystyle f^{-1}(x)=\frac{x-8}{4}}\) (solved for \(f^{-1}(x)\)) \(\\[1.1em]\) then, plug in x=4. \(\\[1.1em]\) \(\color{blue}{\displaystyle f^{-1}(8)=\frac{8-8}{4}=0}\) -------------------------------------- Shortcut. \(\color{blue}{\displaystyle f(x)=4x+8}\) \(\color{blue}{\displaystyle 8=4x+8}\) \(\color{blue}{\displaystyle 0=4x}\) \(\color{blue}{\displaystyle 0=x}\) So, \(\color{blue}{\displaystyle f^{-1}(8)=0}\).

OpenStudy (benlindquist):

so basically do that same thing but, plug in the different equation?

OpenStudy (solomonzelman):

Yes, when you plug in f(x)=4, the solution for x is f^(-1)(4).

OpenStudy (benlindquist):

http://prntscr.com/d351fp it looks like that not just an equation to fill out

OpenStudy (solomonzelman):

yes, then do what I called the official approach.

OpenStudy (benlindquist):

ok

OpenStudy (solomonzelman):

First solve for \(f^{-1}(x)\) to find the inverse function. Swipe x and y, and solve for y. (Your y is now f^(-1)(x).) The, plug in x=4, to find \(f^{-1}(4)\).

OpenStudy (benlindquist):

ok

OpenStudy (solomonzelman):

What do the last two blanks on the first line correspond to?

OpenStudy (benlindquist):

ok

OpenStudy (solomonzelman):

Also, I want to add a brief note about the difference in methods for finding f^(-1)(4). In the "official approach" you are swiping f(x) and x, solve for x (which used to be f(x)), and your f(x) (which used to be x) is (now) an inverse function. Then, to find f^(-1)(4) you plug in x=4. So, (if you just needed to find f^(-1)(4)) you can observe that plugging in f(x)=4 (without finding the inverse function), is precisely the same thing (but faster).

OpenStudy (solomonzelman):

but since you do need \(f^{-1}(x)\), we will find the inverse function first.

OpenStudy (solomonzelman):

Do you know how to find the inverse function in general?

OpenStudy (benlindquist):

yes

OpenStudy (solomonzelman):

Ok, we will start to apply the same steps to \(f(x)=\frac{1}{2}x+4\).

OpenStudy (benlindquist):

so basically i multiply 2 x 1/2 to get rid of the fraction?

OpenStudy (benlindquist):

f(x) + 4?

OpenStudy (solomonzelman):

Well, first (just as matter of convenience, and thus it is typically taught), you write y instead of f(x). So, \(y=\frac{1}{2}x+4\). Then, you swipe \(x\) and \(y\). (Like this) \(x=\frac{1}{2}y+4\). Can you solve for \(y\) now?

OpenStudy (benlindquist):

so then I would use the inverse function?

OpenStudy (solomonzelman):

yes, but first you got to find this inverse function.

OpenStudy (solomonzelman):

Can you solve this for \(y\) ? \(x=\frac{1}{2}y+4\).

OpenStudy (benlindquist):

So x = 1/2y + 4. 2x = y + 2 then subtract - from 2x and 2 -2 and then multiply by 1

OpenStudy (solomonzelman):

can you say that again (rephrase) ?

OpenStudy (solomonzelman):

2x=y+2 is correct, I just don't get what exactly you suggest afterwords.

OpenStudy (benlindquist):

x = 1/2y + 4 2x = y + 2 -2 -2 —————- 2x - 2 = y 2x + 2 = y y = 2x + 2

OpenStudy (solomonzelman):

why does y=2x-2 turn into y=2x+2 ?

OpenStudy (benlindquist):

is it not supposed to become a positive sign?

OpenStudy (solomonzelman):

It is not supposed to. You have y=2x-2 after you subtract (which is correct). The sign doesn't just all of the sudden, for no reason turn positive if it is negative.

OpenStudy (benlindquist):

okk

OpenStudy (solomonzelman):

Ok, so this is the inverse function!

OpenStudy (solomonzelman):

\(f^{-1}(x)=2x-2\)

OpenStudy (solomonzelman):

Alrighty so far?

OpenStudy (benlindquist):

yes

OpenStudy (solomonzelman):

Ok, now you are asked to find \(f^{-1}(4)\).

OpenStudy (solomonzelman):

So, plug in \(x=4\) into \(f^{-1}(x)=2x-2\) ....

OpenStudy (solomonzelman):

\(\large f^{-1}(4)=2\color{red}{\cdot 4}-2={\tiny~}?\)

OpenStudy (benlindquist):

f -1^(4) = 8 -2 = f-1^(4) = 6

OpenStudy (solomonzelman):

Yes, \(f^{-1}(4)=6\).

OpenStudy (solomonzelman):

I think we are good to go now:)

OpenStudy (benlindquist):

http://prntscr.com/d35f3o so like this @SolomonZelman

OpenStudy (solomonzelman):

you gave me the same picture ... (?)

OpenStudy (benlindquist):

that is my final answer

OpenStudy (benlindquist):

are the answers in the correct boxes

OpenStudy (solomonzelman):

what you posted shows only empty boxes (same as before)

OpenStudy (solomonzelman):

(well, maybe just for me, but I see nothing on there)

OpenStudy (benlindquist):

@SolomonZelman 1. box i put 2 after x i put - and at the last box it put 2. in the fourth box i put 6

OpenStudy (solomonzelman):